# Efficient Epistemic Uncertainty Estimation in Regression Ensemble Models Using Pairwise-Distance Estimators

### v3

@@ Abstract
This work introduces an efficient novel approach for epistemic uncertainty estimation for ensemble models for regression tasks using pairwise-distance estimators (PaiDEs). Utilizing the pairwise-distance between model components, these estimators establish bounds on entropy. We leverage this capability to enhance the performance of Bayesian Active Learning by Disagreement (BALD). Notably, unlike sample-based Monte Carlo estimators, PaiDEs exhibit a remarkable capability to estimate epistemic uncertainty at speeds up to 100 times faster while covering a significantly larger number of inputs at once and demonstrating superior performance in higher dimensions. To validate our approach, we conducted a varied series of regression experiments on commonly used benchmarks: 1D sinusoidal data, Pendulum, Hopper, Ant and Humanoid. For each experimental setting, an active learning framework was applied to demonstrate the advantages of PaiDEs for epistemic uncertainty estimation. We compare our approach to existing active learning methods and find that our approach outperforms on high-dimensional regression tasks.
Machine Learning, ICML
[…]
@@ 1 Introduction
In this paper, we propose Pairwise-Distance Estimators (PaiDEs) as a non-sample based method for estimating epistemic uncertainty in deep ensembles with probabilistic outputs for regression tasks. Epistemic uncertainty, often distinguished from aleatoric uncertainty, reflects a model’s ignorance and can be reduced by increasing the amount of data available (Hora, 1996; Der Kiureghian & Ditlevsen, 2009; Hüllermeier & Waegeman, 2021). The significance of epistemic uncertainty is particularly pronounced in safety-critical systems, where a single erroneous prediction could lead to catastrophic consequences (Morales-Torres et al., 2019). Moreover, leveraging epistemic uncertainty proves beneficial as an acquisition criterion for active learning strategies (Houlsby et al., 2011).
@@
Previously Monte Carlo (MC) methods have been employed for estimating epistemic uncertainty, in regression tasks, due to the absence of closed-form expressions in most modeling scenarios (Depeweg et al., 2018; Berry & Meger, 2023). However, as the output dimension increases, these MC methods require a large number of samples to get accurate estimates. PaiDEs offer one non-sample based alternative for estimating information-based criterion in ensemble regression models with probabilistic outputs (Kolchinsky & Tracey, 2017).
@@
Ensembles can be conceptualized as committees, with each ensemble component serving as a committee member (Rokach, 2010). PaiDEs synthesize the consensus amongst a committee of probabilistic learners by calculating the distributional distance between each pair of committee members. Distributional distance is a measure of similarity between two probability distributions. PaiDEs aggregate the pairwise distributional distances in a way that accurately estimates the differential entropy of the entire ensemble. Assuming that the pairwise distances can be efficiently calculated, PaiDEs provide an efficient way to estimate epistemic uncertainty that is not sample-dependent. Figure 1 illustrates a pair of committee members in agreement with low epistemic uncertainty (i.e., small distributional distance) on the right, and another pair in disagreement with high epistemic uncertainty (i.e., large distributional distance) on the left.
This study showcases the application of PaiDEs in estimating epistemic uncertainty on regression tasks for ensembles with probabilistic outputs. Unlike the well-explored area of estimating uncertainty in classification, regression problems pose unique challenges in this field. Information-theoretic quantities such as entropy are more challenging to estimate for an ensemble of continuous distributions compared to an ensemble of categorical distributions, with an ensemble of categorical distributions creating a mixture of categorical distributions which one can easily calculate the entropy for. The primary ensemble utilized in this study is a collection of Normalizing Flows (NFs), known for their effectiveness in capturing heteroscedastic and complex multi-modal aleatoric uncertainty (Rezende & Mohamed, 2015; Kingma & Dhariwal, 2018). This holds particular significance in the realm of robotic locomotion, where these characteristics are pertinent due to the frequent encounters with nonlinear stochastic dynamics. Moreover, these tasks present an ideal application for active learning methods, as acquiring additional data is often resource-intensive and should therefore be conducted with optimal efficiency. Our contributions are as follows:
@@
  • We introduce the framework for applying PaiDEs to estimate epistemic uncertainty in deep ensembles with probabilistic outputs (Section 5).
@@
  • We extend previous epistemic uncertainty estimation methods to higher dimensional settings, and demonstrate how PaiDEs outperform MC and other active learning benchmarks in the higher dimensional setting with rigorous statistical testing (Section 6).
@@
  • We provide an analysis of the time-saving advantages offered by PaiDEs compared to MC estimators for epistemic uncertainty estimation (Section 6.4).
@@ 2 Problem Statement
Following a supervised learning framework for regression, let $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$ denote a dataset, where $x_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}^{d}$, and our objective is to approximate a complex multi-modal conditional probability $p(y|x)$. Let $f_{\theta}(y,x)$ denote our approximation to the conditional probability density, where $\theta$ is a set of parameters to be learned and is distributed as $p(\theta)$.
Leveraging a learned model, $f_{\theta}(y,x)$, we wish to estimate uncertainty which is typically viewed from a probabilistic perspective (Cover & Thomas, 2006; Hüllermeier & Waegeman, 2021). When capturing uncertainty in supervised learning, one common measure is that of conditional differential entropy,
[…]
Utilizing conditional differential entropy, we can establish an estimate for epistemic uncertainty as introduced by Houlsby et al. (2011), expressed as:
$I(y,\theta|x)$ $=H(y|x)-E_{p(\theta)}\left[H(y|x,\theta)\right],$ (1)
where $I(\cdot)$ refers to mutual information. Equation 1 demonstrates that epistemic uncertainty, $I(y,\theta|x)$, can be represented by the difference between total uncertainty, $H(y|x)$, and aleatoric uncertainty, $E_{p(\theta)}\left[H(y|x,\theta)\right]$. Mutual information measures the information gained about one variable by observing another.
@@
$f_{\theta}(y,x)=\sum_{j=1}^{M}\pi_{j}f_{\theta_{j}}(y,x)\qquad\sum_{j=1}^{M}\pi_{j}=1,$ (2)
where $M$ and $0\leq\pi_{j}\leq 1$ are the number of model components and the component weights, respectively. In order to create an ensemble, one of two ways is typically chosen: randomization (Breiman, 2001) or boosting (Freund & Schapire, 1997). While boosting has led to widely used machine learning methods (Chen & Guestrin, 2016), randomization has been the preferred method in deep learning due to its tractability and ease of implementation (Lakshminarayanan et al., 2017).
Utilizing our ensemble we can estimate epistemic uncertainty in decision-making scenarios such as active learning. When all components produce the same $f_{\theta_{i}}(y,x)$, $I(y,\theta|x)$ is zero, indicating no epistemic uncertainty. Conversely, when the components exhibit varied output distributions, epistemic uncertainty is high. In an active learning framework, one chooses, at each iteration, what data points to add to a training dataset such that the model’s performance improves as much as possible (MacKay, 1992; Settles, 2009). In our context, one chooses the $x$’s that maximize Equation 1 and adds those data points to the training set as in Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011). It’s worth noting that in the realm of continuous outputs as in regression and ensemble models, Equation 1 often lacks a closed-form solution, as the entropy of most mixtures of continuous distributions cannot be expressed in closed form (Kolchinsky & Tracey, 2017). Hence, prior methods have resorted to Monte Carlo (MC) estimators for the estimation of epistemic uncertainty (Depeweg et al., 2018; Postels et al., 2020). One such Monte Carlo method samples $K$ points from the model, $y_{j}\sim f_{\theta}(y,x)$, and then estimates the total uncertainty,
$\hat{H}_{MC}(y|x)=\frac{-1}{K}\sum_{j=1}^{K}\ln f_{\theta}(y_{j},x).$
The MC approach approximates the intractable integral over the continuous domain requiring only point-wise evaluation of the ensemble and its components. However, as the number dimensions increases, MC methods typically require a greater number of samples which creates a greater computational burden (Rubinstein & Glynn, 2009).
@@ 3 Pairwise-Distance Estimators
Unlike MC methods, PaiDEs completely remove this dependence on sampling by leveraging generalized distance functions between model component distributions. They can be applied when estimating the entropy of mixture distributions as long as the pairwise-distances have a closed-form and thus can be utilized to estimate $H(y|x)$ from Equation 1. The derivation and properties of PaiDEs follow from Kolchinsky & Tracey (2017); we are extending the use of PaiDEs to a supervised learning problem and epistemic uncertainty estimation.
@@ 3.1 Properties of Entropy
Now that we have a representation of the joint, following principles of information theory (Cover & Thomas, 2006), we can write its entropy as,
@@
Intuitively, the lower bound can be justified by the fact that conditioning on more variables can only decrease or keep entropy the same. The upper bound follows from Equation 3 and $H(\theta|y,x)\geq 0$.
@@ 3.2 PaiDEs Definition
Let $D(p_{i}\;\|\;p_{j})$ denote a generalized distance function between the probability distributions $p_{i}$ and $p_{j}$, which for our case represent $p_{i}=p(y|x,\theta_{i})$ and $p_{j}=p(y|x,\theta_{j})$, respectively. More specifically, $D$ is referred to as a premetric, $D(p_{i}\;\|\;p_{j})\geq 0$ and $D(p_{i}\;\|\;p_{j})=0$ if $p_{i}=p_{j}$. The distance function need not be symmetric nor obey the triangle inequality. As such, PaiDEs can be defined as,
$\hat{H}_{\rho}(y|x)$ $:=H(y|\theta,x)$
$-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D(p_{i}\;\|\;p_{j})\right)}.$ (5)
PaiDEs have many options for $D(p_{i}\;\|\;p_{j})$ (Kullback-Leibler divergence, Wasserstein distance, Bhattacharyya distance, Chernoff $\alpha$-divergence, Hellinger distance, etc.).
@@
Using the extreme distance functions,
@@
$D_{max}(p_{i}\;\|\;p_{j})$ =$=\begin{cases}0,&\text{if }p_{i}=p_{j},\\ \infty,&o/w,\end{cases}$
@@
Refer to Appendix A for the proof of Theorem 3.1. This provides a general class of estimators but a distance function still needs to be chosen. Certain distance functions improve the bounds in Equation 4.
@@ 3.3 Tighter Bounds for PaiDEs
Let the Chernoff $\alpha$-divergence be defined as (Nielsen, 2011),
[…]
where $\alpha\in\left[0,1\right]$.
@@
When applying Chernoff $\alpha$-divergence as our distance function in Equation 5, we achieve a tighter lower bound than $H(y|\theta,x)$,
$\hat{H}_{C_{\alpha}}(y|x)$ $=H(y|\theta,x)$
$-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-C_{\alpha}(p_{i}\;\|\;p_{j})\right)},$ (6)
@@
Refer to Appendix A for the proof of Corollary 3.2 which follows the steps in Kolchinsky & Tracey (2017). In addition, the tightest lower bound can be shown to be $\alpha=0.5$ for certain situations (Kolchinsky & Tracey, 2017). Note that when $\alpha=0.5$, the Chernoff $\alpha$-divergence is known as the Bhattacharyya distance,
@@
We utilized PaiDEs with the Bhattacharyya distance, $\hat{H}_{Bhatt}(y|x)=\hat{H}_{C_{0.5}}(y|x)$, as one proposed improvement to MC estimators.
@@
Note that the KL divergence does not satisfy the triangle inequality nor is it symmetric, thus it is not metric but does suffice as a generalized distance function.
@@
When applying Kullback-Liebler divergence as our distance function in Equation 5, we achieve a tighter upper bound than $H(y,\theta|x)$,
$\hat{H}_{KL}(y|x)$ $=H(y|\theta,x)$
$-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{KL}(p_{i}\;\|\;p_{j})\right)},$ (9)
@@
Refer to Appendix A for the proof of Corollary 3.3 which follows the steps in Kolchinsky & Tracey (2017). In addition to Bhattacharyaa distance, we applied PaiDEs with KL divergence as another proposed improvement to Monte Carlo estimation.
4 Normalizing Flow Ensembles
In this study, we utilize an ensemble technique named Nflows Base, which has previously shown robust performance in estimating both aleatoric and epistemic uncertainty on simulated robotic datasets by leveraging normalizing flows (NFs) to create ensembles (Berry & Meger, 2023). PaiDEs can be employed with any ensemble possessing probabilistic outputs and closed-form distributional distance between ensemble components.

### v4

# Epistemic Uncertainty Estimation in Regression Ensemble Models with Pairwise Epistemic Estimators

@@ Abstract
This work introduces a novel approach, Pairwise Epistemic Estimators (PairEpEsts), for epistemic uncertainty estimation in ensemble models for regression tasks using pairwise-distance estimators (PaiDEs). By utilizing the pairwise distances between model components, PaiDEs establish bounds on entropy. We leverage this capability to enhance the performance of Bayesian Active Learning by Disagreement (BALD). Notably, unlike sample-based Monte Carlo estimators, PairEpEsts can estimate epistemic uncertainty up to 100 times faster and demonstrate superior performance in higher dimensions. To validate our approach, we conducted a varied series of regression experiments on commonly used benchmarks: 1D sinusoidal data, Pendulum, Hopper, Ant, and Humanoid, demonstrating PairEpEsts’ advantage over baselines in high-dimensional regression active learning.
@@ 1 Introduction
In this paper, we propose Pairwise Epistemic Estimators (PairEpEsts) as a non-sample based method for estimating epistemic uncertainty in deep ensembles with probabilistic outputs for regression tasks. Epistemic uncertainty, often distinguished from aleatoric uncertainty, reflects a model’s ignorance and can be reduced by increasing the amount of data available [26, 16, 29]. The significance of epistemic uncertainty is particularly pronounced in safety-critical systems, where a single erroneous prediction could lead to catastrophic consequences [43]. Moreover, leveraging epistemic uncertainty proves beneficial as an acquisition criterion for active learning strategies [27].
@@
Previously Monte Carlo (MC) methods have been employed for estimating epistemic uncertainty, in regression tasks, due to the absence of closed-form expressions in most modeling scenarios [15, 4]. However, as the output dimension increases, these MC methods require a large number of samples to get accurate estimates. PairEpEsts leverage Pairwise-Distance Estimators (PaiDEs), which provide a non-sample-based alternative for estimating information-theoretic criteria in ensemble regression models with probabilistic outputs [35].
@@
Ensembles can be seen as committees, with each component acting as a member [50]. PairEpEsts synthesize the consensus amongst a committee of probabilistic learners by calculating the distributional distance between each pair of committee members. Aggregating these distances allows accurate estimation of the mutual information between the model output and weight distribution. When pairwise distances are efficiently computable, PairEpEsts offer a fast, sample-free method to estimate epistemic uncertainty. Figure 1 shows committee members in agreement with low uncertainty (small distance) and disagreement with high uncertainty (large distance).
This study demonstrates the use of PairEpEsts to estimate epistemic uncertainty in regression ensembles with probabilistic outputs. Unlike classification, regression poses unique challenges because entropy is harder to estimate for mixtures of continuous distributions than for categorical ones. To address this, we focus on Normalizing Flows (NFs), which can capture heteroscedastic and multimodal aleatoric uncertainty [49, 33]. This capability is particularly relevant for robotic locomotion, where nonlinear stochastic dynamics make data acquisition costly and active learning essential. The proposed framework extends epistemic uncertainty estimation to high-dimensional regression tasks that have been largely underserved in the literature. Our contributions are threefold:
@@
  • We introduce the framework PairEpEsts, which applies PaiDEs to estimate epistemic uncertainty in deep ensembles with probabilistic outputs (Section 4).
@@
  • We extend previous epistemic uncertainty estimation methods to settings with higher-dimensional output [4], and demonstrate how PairEpEsts outperform MC and other active learning baselines in these settings with rigorous statistical testing (Section 5).
@@
  • We provide an analysis of the time-saving advantages offered by PairEpEsts compared to MC estimators for epistemic uncertainty estimation (Section 5.4).
@@ 2 Problem Statement
Following a supervised learning framework for regression, let $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$ denote a dataset, where $x_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}^{d}$, and our objective is to approximate a complex multi-modal conditional probability $p(y|x)$. Let $f_{\theta}(y;x)$ denote our approximation to the conditional probability density, where $\theta$ is a set of parameters to be learned and is distributed as $p(\theta)$.
Leveraging a learned model, $f_{\theta}(y;x)$, we wish to estimate uncertainty which is typically viewed from a probabilistic perspective [14, 29]. When capturing uncertainty in supervised learning, one common measure is that of conditional differential entropy,
$H(y|x)=-\int p(y|x)\ln{p(y|x)}dy.$
Utilizing conditional differential entropy, we can establish an estimate for epistemic uncertainty as introduced by Houlsby et al. [27], expressed as:
$I(y,\theta|x)$ $=H(y|x)-H(y|x,\theta),$ (1)
where $I(\cdot)$ refers to mutual information (MI). Equation 1 demonstrates that epistemic uncertainty, $I(y,\theta|x)$, can be represented by the difference between total uncertainty, $H(y|x)$, and aleatoric uncertainty, $H(y|x,\theta)$. MI measures the information gained about one variable by observing another.
@@
$f_{\theta}(y;x)=\sum_{j=1}^{M}\pi_{j}f_{\theta_{j}}(y;x)\qquad\sum_{j=1}^{M}\pi_{j}=1,$ (2)
where $M$, $0\leq\pi_{j}\leq 1$ and $\theta_{j}$ are the number of model components, the component weights and the parameters for the $j$th component, respectively. In order to create an ensemble, one of two ways is typically chosen: randomization [5] or boosting [20]. While boosting has led to widely used machine learning methods [10], randomization has been the preferred method in deep learning [39].
Utilizing our ensemble we can estimate epistemic uncertainty in decision-making scenarios such as active learning. When all components produce the same $f_{\theta_{i}}(y,x)$, $I(y,\theta|x)$ is zero, indicating no epistemic uncertainty. Conversely, when the components exhibit varied output distributions, epistemic uncertainty is high. In an active learning framework, one chooses, at each iteration, what data points to add to a training dataset such that the model’s performance improves as much as possible [41, 52]. In our context, one chooses the $x$’s that maximize Equation 1 and adds those data points to the training set as in Bayesian Active Learning by Disagreement (BALD) [27]. It’s worth noting that in the realm of continuous outputs as in regression and ensemble models, Equation 1 often lacks a closed-form solution, as the entropy of most mixtures of continuous distributions cannot be expressed in closed form [35]. Hence, prior methods have resorted to MC estimators for the estimation of epistemic uncertainty [15, 48]. One such MC method samples $K$ points from the model, $y_{j}\sim f_{\theta}(y;x)$, and then estimates the total uncertainty,
$\hat{H}_{MC}(y|x)=\frac{-1}{K}\sum_{j=1}^{K}\ln f_{\theta}(y_{j};x).$
The MC approach approximates the intractable integral over the continuous domain requiring only point-wise evaluation of the ensemble and its components. However, as the number dimensions increases, MC methods typically require a greater number of samples which creates a greater computational burden [51]. Note that the aleatoric uncertainty can be analytically computed as the average of the entropies of each ensemble component distribution.
@@ 3 Pairwise-Distance Estimators
Unlike MC methods, PaiDEs eliminate sampling by using generalized distance functions between ensemble components. They estimate the entropy of mixture distributions when pairwise distances have closed-form expressions, enabling estimation of $H(y|x)$ from Equation 1. We extend PaiDEs, from Kolchinsky and Tracey [35], to supervised learning and epistemic uncertainty estimation.
@@ 3.1 Properties of Entropy
Using this representation, following principles of information theory [14], we can write its entropy as,
@@
Intuitively, the lower bound can be justified by the fact that conditioning on more variables can only decrease or keep entropy the same. The upper bound follows from Equation 3, and $H(\theta|x,y)\geq 0$ since $\theta$ is modeled as a discrete random variable (ensemble index), unlike $x$ and $y$ which are continuous.
@@ 3.2 PaiDEs Definition
Let $D_{\rho}(p_{i}\;\|\;p_{j})$ denote a generalized distance function between the probability distributions $p_{i}$ and $p_{j}$, which for our case represent $p_{i}=p(y|x,\theta_{i})$ and $p_{j}=p(y|x,\theta_{j})$, respectively. More specifically, $D$ is referred to as a premetric, $D_{\rho}(p_{i}\;\|\;p_{j})\geq 0$ and $D_{\rho}(p_{i}\;\|\;p_{j})=0$ if $p_{i}=p_{j}$. The distance function need not be symmetric nor obey the triangle inequality. As such, PaiDEs can be defined as,
$\hat{H}_{\rho}(y|x)$ $:=H(y|\theta,x)-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{\rho}(p_{i}\;\|\;p_{j})\right)}.$ (5)
PaiDEs have many options for $D_{\rho}(p_{i}\;\|\;p_{j})$ (Kullback-Leibler divergence, Wasserstein distance, Bhattacharyya distance, Chernoff $\alpha$-divergence, Hellinger distance, etc.).
@@
As from Kolchinsky and Tracey [35], using the extreme distance functions,
@@
$D_{max}(p_{i}\;\|\;p_{j})$ $=\begin{cases}0,&\text{if }p_{i}=p_{j},\\ \infty,&otherwise,\end{cases}$
@@
Refer to Appendix A for the proof of Theorem 3.1. This provides a general class of estimators but a distance function still needs to be chosen.
@@ 3.3 Tighter Bounds for PaiDEs
Let the Chernoff $\alpha$-divergence, where $\alpha\in\left[0,1\right]$, be defined as [44],
$D_{C_{\alpha}}(p_{i}\;\|\;p_{j})=-\ln\int p^{\alpha}(y|x,\theta_{i})p^{1-\alpha}(y|x,\theta_{j})dy.$
@@
As from Kolchinsky and Tracey [35], when applying Chernoff $\alpha$-divergence as our distance function in Equation 5, we achieve a tighter lower bound than $H(y|\theta,x)$ from Equation 4,
@@
Refer to Appendix A for the proof of Corollary 3.2. In addition, the tightest lower bound can be shown to be $\alpha=0.5$ for certain situations [35]. This is known as the Bhattacharyya distance,
@@
Note that the KL divergence is a not metric but does suffice as a generalized distance function.
@@
As from Kolchinsky and Tracey [35], when applying Kullback-Liebler divergence as our distance function in Equation 5, we achieve a tighter upper bound than $H(y,\theta|x)$ from Equation 4,
@@
Refer to Appendix A for the proof of Corollary 3.3.
4 Estimating Epistemic Uncertainty with Pairwise Epistemic Estimators
Our extension of prior methodologies enables the estimation of epistemic uncertainty, without the need for sampling. We illustrate this capability for NFs in the main paper and for Probabilistic Network Ensembles (PNEs) in Appendix E [39]. Note that our proposed estimators are applicable to any ensemble model whose component output distributions have closed-form pairwise distances. Since many models have these properties, we focus on Nflows Base for its expressiveness with multimodal aleatoric uncertainty and on widely used PNEs.
4.1 Pairwise Epistemic Estimators
$\hat{I}_{\rho}(y,\theta)$ $=\hat{H}_{\rho}(y|x)-H(y|x,\theta)=-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{\rho}(p_{i}\;\|\;p_{j})\right)}.$ (9)
PaiDEs estimate epistemic uncertainty using only pairwise component distances, removing reliance on sampling. We propose two estimators:
$\hat{I}_{B}(y,\theta)$ $=-\sum_{i=1}^{M}\pi_{i}\ln{\sum_{j=1}^{M}\pi_{j}\exp\left(-D_{B}(p_{i}\;\|\;p_{j})\right)},$
where $D_{B}(p_{i}\;\|\;p_{j})$ and $D_{KL}(p_{i}\;\|\;p_{j})$ are defined for Gaussians in Appendix B. We refer to $\hat{I}_{B}(y,\theta)$ as PairEpEst-Bhatt and $\hat{I}_{KL}(y,\theta)$ as PairEpEst-KL.